OpenStudy (anonymous):

sec^2 5pi/6 - tan^2 5pi/6 how to solve this? please help! thankyou!

OpenStudy (saifoo.khan):

@amistre64

OpenStudy (saifoo.khan):

@satellite73

OpenStudy (anonymous):

please help.

OpenStudy (callisto):

An identity may help you... \[tan^2x+1 = sec^2x\] \[=>sec^2x - tan^2x= 1\] For your question, sec^2 5pi/6 - tan^2 5pi/6, you can see that x=5pi/6... Can you get the answer??

OpenStudy (anonymous):

sec^2 (5pi/6) - tan^2 (5pi/6) = 1 -> (-2 sqrt of 3/ 3) - (- sqrt of 3 / 3) = - sqrt of 3 / 3 = 1 ( <- this is my answer... is it correct? thankyou!)

OpenStudy (anonymous):

@broderick365 help please? thankyou!

OpenStudy (anonymous):

sorry! is it ( - sqrt of 3/ 3) ^2 = 1 ??? :)

OpenStudy (anonymous):

and my teacher told us to SIMPLIFY all our answers. how am i going to simplify this? should the answer supposed to be 3 = 1 ??? :D

OpenStudy (anonymous):

@hartnn please help? what's the exact answer? thankyou! :)

hartnn (hartnn):

the exact answer as told by callisto and you also is 1 .

OpenStudy (anonymous):

oh. thanks. :)

OpenStudy (callisto):

sec^2 (5pi/6) - tan^2 (5pi/6) = 1 Since the angle is the same, you can use that identity to get the answer - 1.

OpenStudy (anonymous):

@Callisto i thought it was 1? thankyou